# LoL-PIM: Long-Context LLM Decoding with Scalable DRAM-PIM System

### v1

@@ Abstract.
The expansion of large language models (LLMs) with hundreds of billions of parameters presents significant challenges to computational resources, particularly data movement and memory bandwidth. Long-context LLMs, which process sequences of tens of thousands of tokens, further increase the demand on the memory system as the complexity in attention layers and key-value cache sizes is proportional to the context length. Processing-in-Memory (PIM) maximizes memory bandwidth by moving compute to the data and can address the memory bandwidth challenges; however, PIM is not necessarily scalable to accelerate long-context LLM because of limited per-module memory capacity and the inflexibility of fixed-functional unit PIM architecture and static memory management. In this work, we propose LoL-PIM– a multi-node PIM architecture that accelerates long context LLM through hardware-software co-design. In particular, we propose how pipeline parallelism can be exploited across a multi-PIM module while a direct PIM access (DPA) controller (or DMA for PIM) is proposed that enables dynamic PIM memory management and results in efficient PIM utilization across a diverse range of context length. We developed an MLIR-based compiler for LoL-PIM– a commercial PIM-based compiler where the software modifications were implemented and evaluated, while the hardware changes were modeled in the simulator. Our evaluations demonstrate that LoL-PIM significantly improves throughput and reduces latency for long-context LLM inference, outperforming both multi-GPU and GPU-PIM systems (up to 8.54$\times$ and 16.0$\times$ speedup , respectively), thereby enabling more efficient deployment of LLMs in real-world applications.
@@ 2.1. Long-context LLM Decoding
Long-context LLM Inference. Long-context LLMs excel at maintaining coherence over tens of thousands of tokens, significantly enhancing contextual relevance across various tasks – e.g., long document summarization (Zhong et al., 2021) generates cohesive summaries from information spread across different sections of extensive text, and repository-level code analysis (Liu et al., 2023) extends programming assistants’ capabilities to analyze entire codebases comprising thousands of lines. Chain-of-thought (CoT) reasoning also improves answer quality by leveraging multi-step contextual reasoning (Wei et al., 2022; OpenAI, 2024). These advancements rely on inference-time processing of extended contexts, with modern LLMs supporting token lengths ranging from 16K to 200K (Bai et al., 2023; MosaicML, 2023; Claude, 2024; OpenAI, 2024). Table 1 summarizes publicly available long-context LLMs, including the number of layers ($n_{l}$), attention heads ($n_{h}$), and the feature dimension of each head ($d_{h}$), as well as structural variations like group query attention (GQA)(Ainslie et al., 2023) and SwiGLU(Shazeer, 2020). (Implementation details are discussed in Sec.7). These models are trained on long-context datasets, enabling them to perform tasks requiring extended input contexts processing. Table 2 highlights the characteristics of popular long-context benchmarks, which exhibit significant variability in context lengths across requests.
Task QMSum HotpotQA Musique mean $13966$ $13465$ $16362$ std $6182$ $3921$ $1651$ max $30456$ $17674$ $17917$ min $2651$ $1917$ $6820$
@@ 2.2. Multi-node Domain-specific PIM System
DRAM-based Domain-specific PIM. DRAM-based Processing-in-Memory (PIM) technology addresses memory-bound challenges by integrating computational capabilities directly within the memory subsystem (Asifuzzaman et al., 2023; Mutlu et al., 2020; Lee et al., 2021b; Kwon et al., 2023b; Lavenier et al., 2016; Choi et al., 2023; Park et al., 2024; Heo et al., 2024). Fixed functional unit PIM systems are optimized for accelerating LLM inference through GEMV operations. These PIM feature processing units adjacent to DRAM banks for parallel dot-product computations, pre-generated PIM commands for simplified control, and limited conditional logic to minimize area overhead  (Lee et al., 2022; Kwon et al., 2023b; He et al., 2020; Kwon et al., 2022, 2023a). For example, AiMX (Lee et al., 2022; Kwon et al., 2023b; He et al., 2020; Kwon et al., 2022, 2023a) includes a vector multiply-accumulate unit in each DRAM bank for 16-element dot-product computations, a shared 2KB global buffer for input data, and a minimal pair of 2-byte output registers per processing unit to transfer results off-chip for module-level processing (e.g., Softmax) using Extra Processing Unit (Fig. 2(a)). Additionally, all channels share a broadcasting input data path to minimize resource usage.
@@
Multi-node PIM System. Fig. 2 (a) illustrates the LLM inference on a multi-node system. The system consists of a host CPU connected with multiple nodes, each consisting of a compute-intensive xPU (i.e., GPU or NPU) connected with multiple PIM modules that provide extensive internal memory bandwidth. As motivated by prior works (Park et al., 2024; Heo et al., 2024; Seo et al., 2024), xPU takes responsibility for the compute-intensive context summarization phase (prefill), while PIM modules handle the memory-bandwidth/capacity-demanding token generation phase (decoding). PIM Controller supports program execution for LLM inference compiled for model-parallel serving. Tensor-parallel (TP) processing (Shoeybi et al., 2019) performs head-wise partitionining of MHA and FFN computations to minimize synchronization, while pipelined parallelism (PP) (Huang et al., 2019; Ma et al., 2024) groups the layers and assigns a node to each layer-group for batch-wise parallel processing. As an example, Fig. 2(b) illustrates a parallel execution of four inference requests in a batch are partitioned over four PIM modules (i.e., $(TP,PP)=(2,2)$), being dynamically managed as the request (= Request1) with the end-of-sequence (EOS) token is replaced with a new request (= Request5).
Execution Flow. Executing LLM inference on PIM hardware involves generating PIM commands (Table 3) to manage tasks such as matrix multiplications and attention mechanisms. Fig. 2(c) illustrates the execution flow, where the PIM module interacts with DRAM to store KV caches and directly compute dot-products. These PIM commands incorporate key configuration details such as layers, attention heads, and batch size. However, since the number of tokens during inference is unknown at compile time, PIM commands must be pre-generated for all potential token ranges up to the maximum size, and memory must be pre-allocated accordingly.
@@ 3. Motivation
3.1. Characteristics of Long-context LLM
@@
  • KV-cache Memory Dominance: The KV-cache consumes a far greater portion of memory compared to weight parameters, as its size is proportional to the context length. Each token generation requires a prefilled KV-cache for Attention computations, leading to significant memory usage as the context length increases (Fig. 3(a)).
@@ 3.2. Challenges of PIM Scaling
Third, the KV-cache’s structure suffers from distortion due to low $d_{h}$, while the aspect ratio of weight parameters becomes increasingly biased as more PIM modules partition them. This biased matrix aspect ratio limits input and output reuse. Fixed-functional PIMs exacerbate this by constraining data transfer paths, resulting in significant I/O overhead. To mitigate this, we propose an I/O-aware buffering mechanism that overlaps data transfer with PIM computation. The resource overhead is minimal, as channels or banks share the buffers. Fig. 4(c) shows a substantial reduction in I/O overhead across varying context lengths (Sec. 6).
@@ 4.1. PIM-aware Partitioning
LLM decoding workloads rely on weight parameter matrices and the KV-cache. Workload partitioning ensures that the dimensions of the matrix are distributed across PIM modules, with each subsection fitting within the memory capacity of the module: $N_{ch}\cdot N_{bank}\times D_{row}\cdot N_{row}$. When partitioning LLM workloads across multiple PIM modules, two primary parallelization methods are used: Pipeline Parallelization (PP) (Huang et al., 2019; Ma et al., 2024), which groups LLM layers into $PP$ pipeline stages, and Tensor Parallelization (TP) (Shoeybi et al., 2019), which divides matrix dimensions within each group by $TP$. While both approaches are widely applied in LLM inference systems, TP is more commonly preferred in multi-node PIM systems. A widely used strategy in prior works, called head-first allocation (HFA) (Heo et al., 2024; Seo et al., 2024; Park et al., 2024), is summarized below (Fig. 5(a)):
@@ 5.1. Managing KV-cache in PIM
PIM accelerates Attention computations by managing the KV-cache within its memory modules. Recent studies (Park et al., 2024; Heo et al., 2024; Seo et al., 2024) have used static memory allocation for the KV-cache, based on the maximum token length, and pre-generated PIM commands with fixed physical addresses (e.g., DOT-PROD(row,col) in Table 3). However, PIM lacks flow control instructions, requiring the compiler to pre-generate commands for all possible token lengths and select the appropriate one at runtime. This approach results in inefficient static memory allocation and token-length-agnostic command generation. Pre-generated commands with fixed operand indices cannot adapt to varying KV-cache sizes encountered during LLM decoding. To address these issues, we make two key observations:
@@
  • Token-length-dependent Operands: Operand indices for PIM commands are determined solely by token length. For example, DOT-PROD operand indices (row) are calculated as $row=T_{cur}/256$, derived from matrix partitioning. The Va2Pa table provides a one-to-one mapping between virtual addresses (Va) and physical addresses (Pa), enabling dynamic identification of the KV-cache data address. Virtual addresses (Va) are assigned logically and continuously, starting from 0 and incrementing by 1 for each data entry. For instance, in Request 1, Va=0 maps to Pa=22, while in Request 2, Va=0 maps to Pa=33. By using the physical address from this mapping, the KV-cache data address (e.g., row=22 or row=33) can be efficiently identified on-the-fly.
@@ 5.2. Direct PIM Access (DPA) Commands
Table 4 outlines the DPA commands designed to enhance the flexibility of PIM commands. These commands encapsulate repetitive token operations, enabling runtime execution without requiring re-compilation. The Dyn-Loop command encodes a loop structure, with Loop-Bound (LB) specifying the number of repetitions and Loop-Entry (LE) defining the number of commands to repeat. These parameters are determined dynamically based on the token length of each request. The Dyn-Modi command adjusts target operand fields (e.g., Target=row for DOT-PROD) of the subsequent PIM command using a coefficient as the stride (e.g., Coef=1 for row-wise traversal of DOT-PROD). These modified operands function as virtual addresses, initialized at zero and incremented by the stride during runtime dispatch. The request ID and virtual address are then used to retrieve the corresponding physical address from the virtual-to-physical memory table. For example, as shown in Fig. 6(c), for Request-2 with $T_{cur}=300$, LB=2 and LE=4. The first DOT-PROD command maps to the physical address (row=33), and the second maps to (row=34). This demonstrates how PIM commands dynamically determine operand physical addresses during runtime execution.
@@ 5.4. Use Case: Lazy Memory Allocation
Integrating the proposed DPA and on-module dispatcher enables lazy memory allocation, which dynamically adjusts PIM memory capacity based on requests’ context lengths. The KV-cache increases with each iteration and eventually exceeds the pre-allocated chunk capacity. At this point, the Va2Pa table is updated to allocate a new chunk to accommodate the additional KV-cache. As shown in Fig. 6(d), when Request-1 requires more capacity, the host allocates a new chunk and updates the Va2Pa table. Notably, the new chunk does not need to be adjacent to the previous one, allowing on-demand allocation in non-contiguous memory regions and improving memory utilization.
Lazy memory allocation also increases the batch size for pipeline parallelization. The host determines the batch size for each iteration of pipeline processing. Under static memory allocation, requests must be assigned memory sufficient for the maximum context length, which limits the number of requests that can be processed simultaneously, even when current requests have smaller context lengths. In contrast, lazy memory allocation uses a smaller number of non-contiguous chunks to handle requests with small KV-cache sizes, significantly boosting average batch size. For example, the average batch size increases by up to 380% compared to static allocation, approaching near-optimal levels (Fig. 4). As discussed in Sec. 4, a larger effective batch size improves PIM’s capacity utilization, reduces pipeline bubbles, and enhances throughput.
@@ 6. PIM I/O Bottleneck
I/O Bottleneck Analysis. While our mapping strategy improves LoL-PIM’s efficiency for LLM inference, it doesn’t address the I/O data transfer overhead inherent to it. LoL-PIM’s buffers are significantly smaller than those in recent DRAM-PIM architectures (Heo et al., 2024; Park et al., 2024), limiting input and output data reuse. Fig. 7(a) reveals that data transfer latency for Out-Reg (DT-Out) and GB (DT-GB) accounts for over 50% of total latency in $QK^{T}$ and $SV$ operations. This highlights the need for architectural improvements in LoL-PIM to reduce input and output data transfer overhead.
@@ 7.1. Multi-PIM Compiler and Runtime
PIM Command Generation. We implemented PIM Dialect in the Codegen Pass to convert the lowered Transformer operations into the PIM command stack. PIM Dialect consists of two main steps: 1) PIM-aware multi-module partitioning and 2) microcoding for PIM command generation. For instance, given the number of PIM modules and workload information (LLM structure, context length, batch size), $QK^{T}$’s GEMV workload is partitioned and distributed to multiple PIM modules (①). For each module, the loop-based PIM command generation step constructs a command stack for the partitioned $QK^{T}$ computation . Since the same computation patterns repeat across Transformer decoder layers and tokens, a PIM command stack is encoded with DPA for compact representation(②). The necessary metadata for runtime dispatching, especially the virtual memory allocation for the weights and KV-cache are also encoded (③).
@@
AiM Memory Configuration GDDR6 16Gb/s; x16 organization;

### v2

@@ Abstract.
The expansion of large language models (LLMs) with hundreds of billions of parameters presents significant challenges to computational resources, particularly data movement and memory bandwidth. Long-context LLMs, which process sequences of tens of thousands of tokens, further increase the demand on the memory system as the complexity in attention layers and key-value cache sizes is proportional to the context length. Processing-in-Memory (PIM) maximizes memory bandwidth by moving compute to the data and can address the memory bandwidth challenges; however, PIM is not necessarily scalable to accelerate long-context LLM because of limited per-module memory capacity and the inflexibility of fixed-functional unit PIM architecture and static memory management. In this work, we propose LoL-PIM– a multi-node PIM architecture that accelerates long context LLM through hardware-software co-design. In particular, we propose how pipeline parallelism can be exploited across a multi-PIM module while a direct PIM access (DPA) controller (or DMA for PIM) is proposed that enables dynamic PIM memory management and results in efficient PIM utilization across a diverse range of context length. We developed an MLIR-based compiler for LoL-PIM– a commercial PIM-based compiler where the software modifications were implemented and evaluated, while the hardware changes were modeled in the simulator. Our evaluations demonstrate that LoL-PIM significantly improves throughput and reduces latency for long-context LLM inference, outperforming both standalone and heterogeneous systems (up to 8.54$\times$ and 16.0$\times$ speedup, respectively), thereby enabling more efficient deployment of LLMs in real-world applications.
@@ 2.1. Long-context LLM Decoding
Long-context LLM Inference. Long-context LLMs excel at maintaining coherence over tens of thousands of tokens, significantly enhancing contextual relevance across various tasks – e.g., long document summarization (Zhong et al., 2021) generates cohesive summaries from information spread across different sections of extensive text, and repository-level code analysis (Liu et al., 2023) extends programming assistants’ capabilities to analyze entire codebases comprising thousands of lines. Chain-of-thought (CoT) reasoning also improves answer quality by leveraging multistep contextual reasoning (Wei et al., 2022; OpenAI, 2024). These advancements rely on inference-time processing of extended contexts, with modern LLMs supporting token lengths ranging from 16K to 200K (Bai et al., 2023; MosaicML, 2023; Claude, 2024; OpenAI, 2024). Table 1 summarizes publicly available long-context LLMs, including the number of layers ($n_{l}$), attention heads ($n_{h}$), and the feature dimension of each head ($d_{h}$), as well as structural variations like group query attention (GQA)(Ainslie et al., 2023) and SwiGLU(Shazeer, 2020). (implementation details are discussed in Sec.7). These models are trained on long-context datasets, enabling them to perform tasks requiring extended input contexts processing. Table 2 highlights the characteristics of popular long-context benchmarks, which exhibit significant variability in context lengths across requests.
@@ 2.2. Multi-node Domain-specific PIM System
DRAM-based Domain-specific PIM. DRAM-based Processing-in-Memory (PIM) technology addresses memory-bound challenges by integrating computational capabilities directly within the memory subsystem (Asifuzzaman et al., 2023; Mutlu et al., 2020; Lee et al., 2021b; Kwon et al., 2023b; Lavenier et al., 2016; Choi et al., 2023; Park et al., 2024; Heo et al., 2024). Fixed functional unit PIM systems are optimized for accelerating LLM inference through GEMV operations. These PIM feature processing units adjacent to DRAM banks for parallel dot-product computations, pre-generated PIM commands for simplified control, and limited conditional logic to minimize area overhead  (Lee et al., 2022; Kwon et al., 2023b; He et al., 2020; Kwon et al., 2022, 2023a). For example (Fig. 2(a)), AiMX (Lee et al., 2022; Kwon et al., 2023b; He et al., 2020; Kwon et al., 2022, 2023a) includes a vector multiply-accumulate unit in each DRAM bank for 16-element dot-product computations, a shared 2KB global buffer for input data, and a minimal pair of 2-byte output registers per processing unit to transfer results off-chip for module-level processing (e.g., Softmax) using Extra Processing Unit. Additionally, all channels share a broadcasting input data path to minimize resource usage.
@@
Multi-node PIM System. Fig. 2 (a) illustrates the LLM inference on a multi-node system. The system consists of a host CPU connected with multiple nodes, each consisting of a compute-intensive xPU (i.e., GPU or NPU) connected with multiple PIM modules that provide extensive internal memory bandwidth. As motivated by prior works (Park et al., 2024; Heo et al., 2024; Seo et al., 2024), xPU takes responsibility for the compute-intensive context summarization phase (prefill), while PIM modules handle the memory-bandwidth/capacity-demanding token generation phase (decoding). PIM Controller supports program execution for LLM inference compiled for model-parallel serving. Tensor-parallel (TP) processing (Shoeybi et al., 2019) performs head-wise partitionining of MHA and FFN computations to minimize synchronization, while pipelined parallelism (PP(Huang et al., 2019; Ma et al., 2024) groups the layers and assigns a node to each layer-group for batch-wise parallel processing. As an example, Fig. 2(b) illustrates a parallel execution of four inference requests in a batch are partitioned over four PIM modules (i.e., $(TP,PP)=(2,2)$), being dynamically managed as the request (= Request2) with the end-of-sequence (EOS) token is replaced with a new request (= Request5).
Execution Flow. Executing LLM inference on PIM hardware involves generating PIM commands (Table 3) to manage tasks such as matrix multiplication and attention mechanism. Fig. 2(c) illustrates the execution flow, where the PIM controller interacts with the PIM channels to store KV caches and directly compute dot products. These PIM commands incorporate necessary LLM configuration details, such as layers, attention heads, and batch size. However, since the number of tokens during inference is unknown at compile time, PIM commands must be pre-generated for all potential token ranges up to the maximum size, and memory must be pre-allocated accordingly.
@@ 3. Motivation
3.1. Characteristics of Long-context LLM Decoding
@@
  • KV-cache Memory Dominance: The KV-cache consumes a far greater portion of memory compared to weight parameters, as its size is proportional to the context length. Each token generation requires a prefilled KV-cache for Attention computations, leading to significant memory usage as the context length increases while FC memory remains the same (Fig. 3(a)).
@@ 3.2. Challenges of PIM Scaling
Third, the KV-cache’s structure suffers from distortion due to low $d_{h}$, since the aspect ratio of weight parameters becomes increasingly biased as more PIM modules partition them. This biased matrix aspect ratio limits input and output reuse. Fixed-functional PIMs exacerbate this by constraining data transfer paths, resulting in significant I/O overhead. To mitigate this, we propose an I/O-aware buffering mechanism that overlaps data transfer with PIM computation. The resource overhead is minimal, as channels or banks share the buffers. Fig. 4(c) shows a substantial reduction in I/O overhead across varying context lengths (Sec. 6).
@@ 4.1. PIM-aware Partitioning
LLM decoding workloads rely on weight parameter matrices and the KV-cache. Workload partitioning (Fig. 1) ensures that the dimensions of the matrix are distributed across PIM modules, with each subsection fitting within the memory capacity of the module: $N_{ch}\cdot N_{bank}\times D_{row}\cdot N_{row}$. When partitioning LLM workloads across multiple PIM modules, two primary parallelization methods are used: Pipeline Parallelization (PP) (Huang et al., 2019; Ma et al., 2024), which groups LLM layers into $PP$ pipeline stages, and Tensor Parallelization (TP) (Shoeybi et al., 2019), which divides matrix dimensions within each group by $TP$. While both approaches are widely applied in LLM inference systems, TP is more commonly preferred in multi-node PIM systems. A widely used strategy in prior works, called head-first allocation (HFA) (Heo et al., 2024; Seo et al., 2024; Park et al., 2024), is summarized below (Fig. 5(a)):
@@ 5.1. Managing KV-cache in PIM
PIM accelerates Attention computations by managing the KV-cache within its memory modules. Recent studies (Park et al., 2024; Heo et al., 2024; Seo et al., 2024) have used static memory allocation for the KV-cache, based on the maximum token length, and pre-generated PIM commands with fixed physical addresses (e.g., DOT-PROD(row,col) in Table 3). However, PIM lacks flow control instructions, requiring the compiler to pre-generate commands for all possible token lengths and select the appropriate one at runtime. This approach results in inefficient static memory allocation and token-length-agnostic command generation. Pre-generated commands with fixed operand indices cannot adapt to varying KV-cache sizes encountered during LLM decoding. To address these issues, we make two key observations of LLM decoding workload:
@@
  • Token-length-dependent Operands: Operand indices for PIM commands are determined solely by token length. For example, DOT-PROD operand indices (row) are calculated as $row=T_{cur}/256$, derived from matrix partitioning. Therefore, if we encode VA as loop dependent index(a operand of command), the Va2Pa table provides a one-to-one mapping between virtual addresses (Va) and physical addresses (Pa), enabling dynamic identification of the KV-cache data address. Virtual addresses (Va) are assigned logically and continuously, starting from 0 and incrementing by 1 for each data entry. For instance, in Request 1, Va=0 maps to Pa=22, while in Request 2, Va=0 maps to Pa=33. By using the physical address from this mapping, the KV-cache data address (e.g., row=22 or row=33) can be efficiently identified on-the-fly.
@@ 5.2. Direct PIM Access (DPA) Commands
Table 4 outlines the DPA commands designed to enhance the flexibility of PIM commands. These commands encapsulate repetitive token operations, enabling runtime execution without requiring re-compilation. The Dyn-Loop command encodes a loop structure, with Loop-Bound (LB) specifying the number of repetitions and Loop-Entry (LE) defining the number of commands to repeat. These parameters are determined dynamically based on the token length of each request. The Dyn-Modi command adjusts target operand fields (e.g., Target=row for DOT-PROD) of the subsequent PIM command using a coefficient as the stride (e.g., Coef=1 for row-wise traversal of DOT-PROD). These modified operands function as virtual addresses, initialized at zero and incremented by the stride during runtime dispatch. The request ID and virtual address are then used to retrieve the corresponding physical address from the virtual-to-physical memory table. For example, as shown in Fig. 6(b), for Request-2 with $T_{cur}=300$, LB=2 and LE=4, the first DOT-PROD command maps to the physical address (row=33), and the second maps to (row=34). This demonstrates how PIM commands dynamically determine operand physical addresses during runtime execution.
@@ 5.4. Use Case: Lazy Memory Allocation
Integrating the proposed DPA and on-module dispatcher enables lazy memory allocation, which dynamically adjusts PIM memory capacity based on requests’ context lengths. The KV-cache increases with each iteration and eventually exceeds the pre-allocated chunk capacity. At this point, the Va2Pa table is updated to allocate a new chunk to accommodate the additional KV-cache. As shown in Fig. 6(d), when Request-1 requires more capacity, the host allocates a new chunk and updates the Va2Pa table. Notably, the new chunk does not need to be adjacent to the previous one in PA, allowing on-demand allocation in non-contiguous memory regions and improving memory utilization.
Lazy memory allocation also increases the effective batch size for pipeline parallelization. The host determines the batch size for each iteration of pipeline processing. Under static memory allocation, requests must be assigned memory sufficient for the maximum context length, which limits the number of requests that can be processed simultaneously, even when current requests have smaller context lengths. In contrast, lazy memory allocation uses a smaller number of non-contiguous chunks to handle requests with small KV-cache sizes, significantly boosting average batch size. For example, the average batch size increases by up to 380% compared to static allocation, approaching near-optimal levels (Fig. 4). As discussed in Sec. 4, a larger effective batch size improves PIM’s capacity utilization, reduces pipeline bubbles, and enhances throughput.
@@ 6. PIM I/O Bottleneck
I/O Bottleneck Analysis. While our mapping strategy improves LoL-PIM’s efficiency for LLM inference, it does not address the I/O data transfer overhead inherent to it. LoL-PIM’s buffers are significantly smaller than those in recent DRAM-PIM architectures (Heo et al., 2024; Park et al., 2024), limiting input and output data reuse. Fig. 7(a) reveals that data transfer latency for Out-Reg (DT-Out) and GB (DT-GB) accounts for over 50% of total latency in $QK^{T}$ and $SV$ operations. This highlights the need for architectural improvements in LoL-PIM to reduce input and output data transfer overhead.
@@ 7.1. Multi-PIM Compiler and Runtime
PIM Command Generation. We implemented PIM Dialect in the Codegen Pass to convert the lowered Transformer operations into the PIM command stack. As shown in Fig. 8 PIM Dialect consists of two main steps: 1) PIM-aware multi-module partitioning and 2) microcoding for PIM command generation. For instance, given the number of PIM modules and workload information (LLM structure, context length, batch size), $QK^{T}$’s GEMV workload is partitioned and distributed to multiple PIM modules (①). For each module, the loop-based PIM command generation step constructs a command stack for the partitioned $QK^{T}$ computation . Since the same computation patterns repeat across Transformer decoder layers and tokens, a PIM command stack is encoded with DPA for compact representation(②). The necessary metadata for runtime dispatching, especially the virtual memory allocation for the weights and KV-cache are also encoded (③).
@@
AiM(Kwon et al., 2023c; Yongkee Kwon et al., 2024) Memory Configuration GDDR6 16Gb/s; x16 organization;